A transformer has 100 loops on the primary coil and 1000 loops on the secondary coil. If the primary coil is connected to 120 VAC, what will the voltage be between the leads in the secondary coil?
Answer #1

We know

\frac{\text{Voltage in Primary coil}}{\text{Voltage in Secondary coil}}=\frac{\text{No of turns in Primary coil}}{\text{No of turns in Secondary coil}}\\ \implies \frac{120}{V_S}=\frac{100}{1000}\\ \implies 1200V=V_S

It is a step up transformer